# FG-PINNs: A neural network method for solving nonhomogeneous PDEs with high frequency components

### v1

@@ 3. Frequency-guided physics-informed neural networks
Figure 7 (B) and (C) depict the temperature distribution and prediction at the initial time, respectively. The numerical results show that FG-PINNs provide an accurate prediction. Figure 7 (A) shows the corresponding absolute error, where the maximum absolute error and the relative $L^{2}$ error are $1.77\times 10^{-2}$ and $6.21\times 10^{-3}$, respectively. The numerical results indicate that FG-PINNs effectively reconstruct the initial temperature distribution. Figure 8 (A) and (B) display the exact solution (t=0.5) and the prediction (t=0.5) of FG-PINNs, respectively. The absolute error is shown in Figure 8 (C), with maximum absolute and relative $L^{2}$ errors of $2.98\times 10^{-3}$ and $1.53\times 10^{-3}$, respectively. Figure 8 (D) illustrates the relative $L^{2}$ error curve. It can be observed that the relative $L^{2}$ error curve oscillates significantly during the training process, which may be attributed to the ill-conditioned nature of the problem itself. After 6,000 iterations, FG-PINNs yield an accurate approximation.
@@
Table 7. Relative $L^{2}$ error at different numbers of modules and training points in Example 5
@@
Table 8. Relative $L^{2}$ error for different activation functions in Example 5
   Activation function    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ error    $1.53\times 10^{-3}$    $2.13\times 10^{-3}$    $2.95\times 10^{-3}$    $6.13\times 10^{-3}$
@@ 5. Summary and discussion
In this work, we developed frequency-guided physics-informed neural networks (FG-PINNs) for solving nonhomogeneous PDEs. FG-PINNs consist of two subnetworks: a high-frequency subnetwork and a low-frequency subnetwork. We employ the high-frequency subnetwork, incorporating the source term, to capture the high-frequency components of the target solution. Inspired by spectral bias, we utilize a fully connected neural network to capture the low-frequency components of the target solution. To evaluate the performance of FG-PINNs, we conducted a series of tests. The numerical results demonstrate that FG-PINNs can rapidly capture both the high-frequency and low-frequency components of the target solution.
Compared to existing methods that require multiple subnetworks, FG-PINNs only need two subnetworks, thereby saving the associated computational overhead. Meanwhile, FG-PINNs exhibit extremely fast convergence speed, as evidenced by the aforementioned series of experiments. This accelerated convergence stems from the fact that FG-PINNs inherently incorporate the frequencies of the target solution. Consequently, both the high-frequency and low-frequency components of the target solution achieve rapid convergence when using FG-PINNs to solve nonhomogeneous PDEs. In contrast, existing methods typically require tens or hundreds of thousands of iterations to obtain an accurate approximation. However, it is important to note that FG-PINNs are currently only applicable to nonhomogeneous problems. Extending this method to homogeneous problems is one of our future research goals.

### v2

@@ 3. Frequency-guided physics-informed neural networks
Figure 7 (B) and (C) depict the temperature distribution and prediction at the initial time, respectively. The numerical results show that FG-PINNs provide an accurate prediction. Figure 7 (A) shows the corresponding absolute error, where the maximum absolute error and the relative $L^{2}$ error are $1.77\times 10^{-2}$ and $6.21\times 10^{-3}$, respectively. The numerical results indicate that FG-PINNs effectively reconstruct the initial temperature distribution. Figure 8 (A) and (B) display the exact solution (t=0.5) and the prediction (t=0.5) of FG-PINNs, respectively. The absolute error is shown in Figure 8 (C), with maximum absolute error and relative $L^{2}$ error of $2.98\times 10^{-3}$ and $1.53\times 10^{-3}$, respectively. Figure 8 (D) illustrates the relative $L^{2}$ error curve. It can be observed that the relative $L^{2}$ error curve oscillates significantly during the training process, which may be attributed to the ill-conditioned nature of the problem itself. After 6,000 iterations, FG-PINNs yield an accurate approximation.
@@
Table 7. Relative $L^{2}$ errors at different numbers of modules and training points in Example 5
@@
Table 8. Relative $L^{2}$ errors for different activation functions in Example 5
   Activation functions    sin(x)    cos(x)    tanh(x)    sigmoid(x)
   Relatvie $L^{2}$ errors    $1.53\times 10^{-3}$    $2.13\times 10^{-3}$    $2.95\times 10^{-3}$    $6.13\times 10^{-3}$
@@ 5. Summary and discussion
In this work, we developed frequency-guided physics-informed neural networks (FG-PINNs) for solving PDEs with high-frequency components. The proposed algorithm relies on prior knowledge about high-frequency components obtained from PDEs. It aims to use this prior knowledge to guide neural networks in rapidly approximating the solution’s high-frequency components. FG-PINNs consist of two subnetworks: a high-frequency subnetwork and a low-frequency subnetwork. We employ the high-frequency subnetwork, incorporating the source terms or initial/boundary conditions with high-frequency components, to capture the high-frequency components of the solution. Inspired by spectral bias, we utilize a fully connected neural network to capture the low-frequency components of the solution. To evaluate the performance of FG-PINNs, we conducted a series of tests. The numerical results demonstrate that FG-PINNs can provide an accurate approximation for PDEs with high-frequency components.
Compared to existing methods that require multiple subnetworks, FG-PINNs only need two subnetworks, thereby saving the associated computational overhead. Meanwhile, FG-PINNs exhibit rapid convergence, as evidenced by the aforementioned series of experiments. This accelerated convergence stems from the fact that FG-PINNs incorporate the high-frequency information of the solution. Therefore, both the high-frequency and low-frequency components of the solution achieve rapid convergence when using FG-PINNs to solve PDEs with high-frequency components. In contrast, existing methods typically require tens of thousands to hundreds of thousands of iterations to obtain an accurate approximation. However, it should be noted that FG-PINNs are extremely dependent on prior information about high-frequency components obtained from PDEs.